OpenStudy (anonymous):

Jill has been watching the squirrels that live in her neighborhood. The number of squirrels changes each week. n f(n) 1 8 2 12 3 18 4 27 Which function best shows the relationship between n and f(n)?

OpenStudy (anonymous):

f(n) = 12(1.5)n f(n) = 12(1.5)n-1 f(n) = 8(1.5)n f(n) = 8(1.5)n-1